find the slope and the y intercept of
y = -5/6 x -6/5

Found 4 solutions by user_dude2008, JATAZAH1000, london maths tutor, Alan3354:
Answer by user_dude2008(1862)

slope: -5/6, y-intercept: (0,-6/5)

Answer by JATAZAH1000(3)
Answer by london maths tutor(243)

This is a linear equation with the general form :
y = mx + c
where m = gradient and c = y intercept
therefore,
y = -5/6 x -6/5
m = - 5/6 and c = -6/5
Answer: gradient = -5/6 and y intercept = -1 1/5

Answer by Alan3354(69443)

It's in the slope-intercept form.
m = -5/6
y-int = (0,-6/5)